OpenStudy (anonymous):

graph the equation by plotting three points -x-y=1

OpenStudy (anonymous):

Rearrange to isolate y. Proceed to plotting.

OpenStudy (anonymous):

how do I do that

OpenStudy (anonymous):

like this y=x-1

OpenStudy (anonymous):

Not quite. When you moved the x over to the other side, what you should have was -y= 1+x To isolate y, you must multiply every term in the equation (term being every chunk separated by + or - from each other) by -1

OpenStudy (anonymous):

multiply by what number? -y

OpenStudy (anonymous):

You must realize that multiplying everything by -y would get you a y in every term of the equation, which is the opposite of what you want to be doing. The goal here is to make y appear only once and without anything else by its side. You should also realize that -y is another way of writing (-1)y.

OpenStudy (anonymous):

What you were doing initially was very close to what you were supposed to do, but you missed the division by -1 on the x term when you isolated the y.

OpenStudy (anonymous):

so dived all terms by -1

OpenStudy (anonymous):

Yes. I said multiply before only because it was -1 and would have been no different from dividing by -1. Sorry if what I said was confusing.

OpenStudy (anonymous):

its fine so y=-1+x

OpenStudy (phi):

you start with -x-y=1 add +x to both sides -y = x + 1 divide all terms by -1 y = -x -1

OpenStudy (anonymous):

Still no. 1 and x are two different terms, you must divide both by -1, not just the 1.

OpenStudy (phi):

now pick small values for x, and find the corresponding y

OpenStudy (anonymous):

0,1,2 for x

OpenStudy (phi):

and y is ?

OpenStudy (anonymous):

(0,-2) (1,-2) (2,-3)

OpenStudy (phi):

typo ?

OpenStudy (anonymous):

is that wrong

OpenStudy (phi):

y= -x - 1 what is y when x=0 ?

OpenStudy (anonymous):

-1

OpenStudy (phi):

yes.

OpenStudy (anonymous):

ok (0,-1) (1,-2) (2,-3)

OpenStudy (anonymous):

is that right

OpenStudy (phi):

yes. you can always check by using the original equation -x-y=1 to test point (2,-3) replace x with 2 and y with -3 : -2 - (-3) = ? 1 -2 +3 =? 1 1 = ? 1 yes. the other points also work

OpenStudy (anonymous):

thanks would you mind helping me with a few more

OpenStudy (phi):

one more